OpenStudy (anonymous):

You just purchased a cellular phone and are trying to determine to which cell phone company you will give your business. When you contacted the Talks-A-Lot Company, they were offering a monthly plan of $40 for 600 minutes and $0.35 for each minute exceeding the 600 minutes. In the Sunday paper you see an ad for the Chat-Away Company, which offers a monthly plan of $50 for 600 minutes and $0.10 for each minute exceeding the 600 minutes. Solve the linear system using the substitution method. Please make sure to solve for both x and y.

OpenStudy (anonymous):

OK. if y is cost and x is minutes, you can represent that by: y=40 + .35(x-600) y=50+10(x-600) (assuming you're solving for the point at which they cost the same, x>600 because if x<600 they're clearly not the same price) Multiply that out and solve.

OpenStudy (anonymous):

ooo i have been doing this wrong. i wrote y =600 + .35

OpenStudy (anonymous):

You need to include the x. Otherwise, you've got this.

OpenStudy (anonymous):

y = 40 + .35x - 210 y - 50 + .10x - 60 ?

OpenStudy (anonymous):

Answer: y=0.35x+40 y=0.10x+50 36 minutes They r equal at 54$ Talks a lot would b better 4 1000

OpenStudy (anonymous):

@rouge lol thanks. but i need to show my work. && i Have no idea how to do the substitution method

OpenStudy (anonymous):

x>600 TalkALot: f(x)=40+0.35(x-600) ChatAway: f(x)=50+0.1(x-600) point of equality: 40+0.35(x-600)=50+0.1(x-600) 0.25(x-600)=10 x-600=40 x=640 y=50+0.1(640-600)=50+4=54 or y=40+0.35(640-600)=40+14=54 you should hire the second company (ChatAway) f(1000)=40+0.35*400 = 180 f(1000)=50+0.1*400 = 90